# Large Language Model-based Data Science Agent: A Survey

### v1

@@ Abstract
11footnotetext: $\dagger$ These authors contributed equally to this work.
@@ 1 Introduction
In recent years, the rapid development of Large Language Models (LLMs) has driven significant innovations across various domains. Leveraging their remarkable capabilities in understanding and generating human-like text, LLMs have become foundational in creating intelligent agents capable of performing complex tasks autonomously. These agents have demonstrated substantial potential in diverse fields, including healthcare Qiu et al. (2024), finance Yu et al. (2024), education Zhang et al. (2025b), and software engineering Hong et al. (2023).
@@
Among these fields, data science has emerged as a particularly critical area for applying LLM-based agents Sun et al. (2024b). Data science involves extracting meaningful insights from vast and diverse datasets, a process that traditionally requires extensive manual effort and expertise. Consequently, LLM-based data science agents (DS Agents) have attracted attention for their ability to automate and optimize data analysis, model development, and decision-making processes.
@@
In this survey, we examine LLM-based data science agents from two complementary perspectives: agent design and data science application. From the agent design perspective, we summarize key architectural paradigms—including single-agent systems, collaborative multi-agent structures, and dynamic agent generation—and analyze core components such as agent roles, execution strategies, knowledge integration, and reflection mechanisms. From the data science perspective, we explore how LLM agents are applied across major workflow stages such as data preprocessing, modeling, evaluation, and visualization. We also outline common task types, including model-building and insight-generation tasks, and characterize the iterative nature of the data science loop. Additionally, our survey goes beyond mere documentation by synthesizing insights from recent studies to identify research opportunities and future directions in this evolving field Sahu et al. (2024); Fan et al. (2023).
From the agent perspective (§3), we summarize how the agent structure is designed (§3.1), how the reasoning of LLM within agents is performed (§3.2), where the knowledge comes from (§3.3), and the reflection of the agent (§3.4). From this perspective, we provided a comprehensive analysis of the agent design of LLM-based data scientist agents.
From the data science perspective (§4), we summarize how LLM-based agents are applied across key stages of the data workflow, including data preprocessing, modeling, evaluation, and visualization. We categorize common task types such as model development and insight generation, and highlight the recurring data science loop that these agents help automate. This perspective clarifies the practical roles LLM agents play in enabling end-to-end data analysis.
@@ 2.1 Modular Architectures
Existing surveys often decompose LLM-based agents into functional modules such as planning, memory, perception, and action. For example, Guo et al. (2024b) identifies components like agent-environment interface and capability acquisition; similar structures appear in Liu et al. (2024a), Sun et al. (2024b), and others.
@@ 2.2 Collaboration and Communication
Agent collaboration is commonly categorized by structure (centralized, decentralized) or mode (cooperation, competition, etc.), as seen in Tran et al. (2025). Related works Guo et al. (2024b), Li et al. (2024g) discuss role-based or layered designs.
@@ 2.3 Capability Evolution and Reflection
Several surveys recognize that agent systems can adapt via feedback, memory updates, or reflection—for example, Li et al. (2024g) includes an evolution phase with memory consolidation, and Wang et al. (2024a) discusses reflective planning. Similar notions appear in Guo et al. (2024b) and others.
@@ 3 Analysis from Agent Perspective
Large Language Model (LLM)-based agents have emerged as powerful tools in various domains, particularly in data science. The design and functionality of LLM-based agents can be understood through their basic components, which include agent role, execution structure, knowledge, and reflection.
Agent role (see §3.1). LLM agents are allocated different roles, which allows the agents to split the main tasks and focus on specific tasks. Diverse agent roles are presented in previous works, ranging from single-agent systems handling all tasks independently to multi-agent systems with specialized roles like developers, testers, planners, etc.
Execution structure (see §3.2). The execution structure designs how agents manage task allocation, task execution, user interaction, error handling, etc. The execution structure covers dynamic planning where agents adjust plans based on real-time feedback, fixed workflows with predefined task sequences, and plan-then-execute frameworks that separate strategy formulation from task execution.
External Knowledge (see §3.3). The knowledge sources allow agents to access and integrate external information, enhancing their ability in specific domains. LLM-based agents augment their knowledge through external databases, retrieval-based approaches, and API calls.
Reflection (see §3.4). Reflection methods provide feedback information for LLM-based agents to improve performance and adapt to complex environments. Techniques include agent feedback for self-correction, model metrics feedback for optimization, code error handling for reliability, and history window mechanisms for long-term learning.
@@ 3.1.1 Single Agent
The single-agent design is the simplest design structure as shown in Figure 3. Most existing single-agent works directly follow the ReAct design (a prompting technique , where LLMs are used to generate both reasoning traces and task-specific actions) Freimanis & Andersson Rhodin (2024); Chen et al. (2024a); Sun et al. (2024a); Jing et al. (2024); Hu et al. (2024a); Deng et al. (2024); Le et al. (2023); Zhang et al. (2024a); Gupta et al. (2024)Bendinelli et al. (2025); Xu et al. (2025), where the single agent will perform thought, action, and observation processes iteratively on its own.
@@ 3.1.2 Two Agents
Planner and executor. Different from the single agent discussed previously, some works split the single agent into two agents Huang et al. (2024b); Liu et al. (2024b); Zhang et al. (2023b); Chi et al. (2024); Li et al. (2024e)Wang et al. (2025a), planner and executor. As an example shown in Figure 4, the planner will get observations from previous execution results or users’ requests, and generate a next-step plan, or a whole plan in advance. Then, the executor is prompted to interact with the environments and will follow the generated plans to execute.
@@
Coder and reviewer. Another type of two-agent structure is the “coder and reviewer” style Trirat et al. (2024); Huang et al. (2023) as shown in Figure 5. In such a design, the coder is responsible for completing tasks following the typical ReAct structure. Another agent, the reviewer is introduced to check the validity of the code generated by the coder. Trirat et al. (2024) allows the reviewer to check the generated response at each step of generation, while Huang et al. (2023) only allows the reviewer to function at the end of the entire generation process.
@@ 3.1.3 Multiple Agents
Software engineering-style team. The Software Engineering (SE) team-style agent design draws inspiration from traditional human software development teams Qian et al. (2024); Zhao et al. (2024b); Trirat et al. (2024); Hong et al. (2023); Tao et al. (2024); Lin et al. (2024); Nguyen et al. (2024).In an SE team-style framework, agents are typically assigned roles that correspond to key human roles in software development, here we state some example roles: The product manager defines the product vision and organizes requirements. The requirements analyst translates user needs into detailed software specifications. The scrum master facilitates task planning and sprint coordination. Hierarchical roles like team leader, module Leader, and function coordinator handle task decomposition at varying levels of granularity. The executor for the task is the developer, who is responsible for implementing the code, while the senior developer refines it. Finally, the QA engineer ensures quality through rigorous testing, and the tester validates specific functionalities. Some SE team-style agents also perform hierarchical roles which allow some agents with high permissions to manage others.
@@
Minimum function agents. Minimum function agents are designed to handle narrowly scoped and atomic tasks separately, as shown in Figure 6. Across existing frameworks, minimum function agents are given very small and specific functions to handle. For instance, code search agents locate relevant files, classes, or methods within a repository, while fault localization agents identify buggy code sections using debugging techniques like spectrum-based fault localization Zhang et al. (2024e). Other agents specialize in generating patches to fix identified issues, executing tests to validate code correctness, or systematically building repository structures from high-level descriptions Zan et al. (2024); Arora et al. (2024). In the context of exception handling, some agents detect fragile code, identify exception types, and implement robust handling mechanisms to enhance code reliability Zhang et al. (2024b). All works emphasize task decomposition and modularity, with outputs from one agent often serving as inputs for another, forming structured and collaborative workflows Zan et al. (2024); Phan et al. (2024); Arora et al. (2024)Seo et al. (2025); You et al. (2025); Ou et al. (2025).
@@
Client-server design. Client-server agents, adopt a hierarchical architecture where a central controller agent manages and coordinates the operations of multiple specialized client agents Zhang et al. (2023a); Yang et al. (2024a); Gandhi et al. (2024); Zhao et al. (2024a). The controller or server agent functions as a project manager, planning entire workflows of the given tasks, and allocates the tasks to the client agents. Client agents perform as roles like software engineers or testers, focusing on executing specific subtasks such as data analysis, modeling, or feedback generation. These roles allow for clear separation of tasks, while also enabling dynamic adjustments based on task requirements or external feedback Zhang et al. (2023a); Bai et al. (2024); Zhao et al. (2024a); Shen et al. (2024).
@@ 3.1.4 Dynamic Agents
Dynamic agents represent a class of multi-agent systems that emphasize adaptability by dynamically creating, modifying, or expanding agents during runtime. Unlike static agents, which follow predefined workflows and configurations, dynamic agents are designed to respond to the complexity or variability of tasks by adjusting the agents’ internal structure (prompt, etc.) or introducing new agents Hu et al. (2024b); Ishibashi & Nishimura (2024). Current frameworks for dynamic agent creation adopt two primary paradigms:
@@
Hierarchical agent generation. This paradigm involves a parent agent or a high-level controller (e.g., Mother Agent in SoA Ishibashi & Nishimura (2024)) that decomposes complex tasks into subtasks and creates child agents to handle the subtasks. Each child agent operates independently on its specific subtask. It is particularly effective in managing tasks with clear functional divisions, such as modular code generation or system-level software design Ishibashi & Nishimura (2024).
@@
Iterative agent generation through feedback. Dynamic agents adjust their structures and behaviors iteratively based on real-time feedback from their environment or other agents (Figure 9). EvoMAC Hu et al. (2024b) exemplifies this paradigm, employing a collaborative rather than hierarchical approach, where agents refine their outputs and workflows through mechanisms analogous to backpropagation—such as a loss agent computing errors and an update agent adjusting agent workflows accordingly. This textual feedback enables the dynamic creation or reconfiguration of agents, supporting adaptability for tasks with evolving requirements, such as changing software specifications (Figure 9) Hu et al. (2024b).
@@ 3.1.5 Summary of Agent Role
    With a manager: In this structure, a central agent manages and controls all agents. Software engineering-style agents and Client-server agents mainly have this structure. For example, in the AutoML-GPTTrirat et al. (2024) framework, a central LLM serves as the controller, managing the entire pipeline by integrating specialized agents for subtasks such as model design and hyperparameter tuning.
@@
    Without a manager: Each agent operates independently and solves tasks autonomously. Minimum function agents mainly pose this structure, since all the agents with minimum function share the same position. An example of this is the MASAI Arora et al. (2024) framework, which utilizes decentralized agents that collaborate on machine learning and data science tasks by sharing results but not a central management system.
@@
    Hierarchical managers: A higher-level agent controls lower-level agents in a layered structure. Hierarchically generated agents and part of software engineering style agents mainly pose such structure. An example of this can be found in Hierarchical agent generation, such as in EvoMAC Hu et al. (2024b), where a parent agent dynamically creates child agents to handle specific subtasks during runtime.
@@
    Compete: Agents work against each other to complete a task, often in the coder-reviewer paradigm (similar to the adversarial concept in GAN). In frameworks like MASAI Arora et al. (2024), agents engage in competitive strategies to address machine learning and data science challenges. The competition between reviewers and coders allows the iterative refinement of the code. Some works also allow multiple agents to propose multiple plans to compete.
@@
    Collaborate: Agents work together toward a shared objective. For example, in the MAGIS Tao et al. (2024) framework, agents assume different roles like Manager, Developer, and QA Engineer to collaborate on resolving GitHub issues. Their tasks are divided to ensure modular development, with continuous collaboration between agents.
@@
    Hybrid: Agents alternate between competing and collaborating based on the task requirements. For instance, in AutoCodeRover Zhang et al. (2024e), agents work together to localize faults and generate patches, but may compete in terms of optimizing solutions or strategies based on the specific issue at hand.
@@
    Static Task Allocation: In some systems, agents are assigned a fixed set of tasks that they perform. For example, in the Data Director Hong et al. (2024) framework, agents follow a static task allocation, where the tasks are predefined and agents work through structured stepwise execution.
@@
    Dynamic Task Allocation: Tasks are allocated based on the real-time needs and feedback from the system or environment. An example of dynamic task allocation can be seen in the EvoMAC Hu et al. (2024b) framework, where agents adjust dynamically based on environmental feedback, creating or dismissing agents as needed to refine or expand their tasks.
@@
    Coarse Granularity: Some agents are given broader, less detailed tasks. For example, in AutoML-GPT Trirat et al. (2024), the central controller agent coordinates the entire machine learning pipeline, handling coarse-grained tasks such as managing the overall workflow rather than focusing on the details of each individual task.
@@
    Fine Granularity: In other cases, tasks are broken down into smaller units for more specific execution. For example, in MapCoder Islam et al. (2024a), multiple agents collaborate, each handling a fine-grained task such as code generation, debugging, or retrieval. This detailed task assignment ensures high accuracy but requires more computation overhead.

### v2

@@ 1 Introduction
In recent years, the rapid development of Large Language Models (LLMs) has driven significant innovations across various domains. Leveraging their remarkable capabilities in understanding and generating human-like text, LLMs have become foundational in creating intelligent agents capable of performing complex tasks autonomously. These agents have demonstrated substantial potential in diverse fields, including healthcare qiu2024llm, finance yu2024fincon, education zhang2025eduplanner, and software engineering hong2023metagpt.
@@
Among these fields, data science has emerged as a particularly critical area for applying LLM-based agents sun2024statistics. Data science involves extracting meaningful insights from vast and diverse datasets, a process that traditionally requires extensive manual effort and expertise. Consequently, LLM-based data science agents (DS Agents) have attracted attention for their ability to automate and optimize data analysis, model development, and decision-making processes.
A central goal of this survey is to clarify how LLM agents can be organized to perform the functions of a data scientist. This leads to a sequence of increasingly detailed research questions. (RQ1) At the organizational level, how should agents’ roles and responsibilities be defined so that complex analytical tasks can be effectively decomposed? (RQ2) Given these roles, how should each agent carry out its reasoning and tool use to ensure reliable execution? (RQ3) Execution usually depends on information beyond the agent’s internal context. What external knowledge should agents access, and how should such knowledge be integrated into their decisions? (RQ4) Even with access to the necessary knowledge, agents still make mistakes, raising the question of what reflection and self-correction mechanisms are required to detect and repair errors during operation. (RQ5) Finally, data science workflows are inherently iterative, prompting the question of how these agent-level capabilities can support end-to-end, loop-based processes rather than isolated single-step tasks.
@@
To address these research questions, we examine LLM-based data science agents from two complementary perspectives: agent design and data science application. From the agent design perspective, we summarize key architectural paradigms—including single-agent systems, collaborative multi-agent structures, and dynamic agent generation—and analyze core components relevant to RQ1–RQ4, including how the agent structure is designed (§3.1), how the reasoning of LLM within agents is performed (§3.2), where the knowledge comes from (§3.3), and the reflection of the agent (§3.4). such as agent roles, execution strategies, knowledge integration, and reflection mechanisms. From the data science perspective(§4), we explore how LLM agents are applied across major workflow stages such as data preprocessing, modeling, evaluation, and visualization. We also outline common task types, including model-building and insight-generation tasks, and characterize the iterative nature of the data science loop. This perspective directly informs RQ5 by illustrating how LLM agents support each stage of an end to end data science workflow. Additionally, our survey goes beyond mere documentation by synthesizing insights from recent studies to identify research opportunities and future directions in this evolving field sahu2024insightbench; fan2023static.
@@ 2.1 Modular Architectures
Existing surveys often decompose LLM-based agents into functional modules such as planning, memory, perception, and action. For example, guo2024llmmultiagentsurvey identifies components like agent-environment interface and capability acquisition; similar structures appear in liu2024llmagentssoftware, sun2024statistics, and others.
@@ 2.2 Collaboration and Communication
Agent collaboration is commonly categorized by structure (centralized, decentralized) or mode (cooperation, competition, etc.), as seen in tran2025multicollaborationllms. Related works guo2024llmmultiagentsurvey, li2024llmmultiagentworkflow discuss role-based or layered designs.
@@ 2.3 Capability Evolution and Reflection
Several surveys recognize that agent systems can adapt via feedback, memory updates, or reflection—for example, li2024llmmultiagentworkflow includes an evolution phase with memory consolidation, and wang2024llmautonomousagents discusses reflective planning. Similar notions appear in guo2024llmmultiagentsurvey and others.
@@ 3 Analysis from Agent Perspective
LLMs provide strong linguistic and generative capabilities, but they exhibit structural limitations that become particularly evident in data science workflows. First, LLMs are prone to hallucination, e.g., fabricating column names, statistical results, or entire processing steps. Second, LLM-generated code is brittle. Even when logically correct in natural language, produced scripts frequently break due to issues such as type inconsistencies, missing dependencies, unstable pipelines, or non-reproducible execution paths. Third, LLMs struggle with long-horizon tasks. Their limited ability to retain and manipulate extended context leads to goal drift, forgotten assumptions, and inconsistencies across multi-step analytical processes. These limitations show that raw LLMs are not reliable as end-to-end data science systems and need structured agent architectures to provide the capabilities they lack and to address common failure modes.
Figure 2: Keyword cloud summarizing the most frequent concepts across the surveyed LLM-based data science agent papers. Larger words represent higher occurrence frequency. The colors in the cloud correspond to the major analysis dimensions discussed in this survey: red words represent agent role design (§3.1), blue words represent execution structure (§3.2), green words represent external knowledge integration (§3.3), and purple words represent reflection mechanisms (§3.4).
Motivated by these limitations, recent work has increasingly turned to agent architectures that introduce structure, modularity, and external support, enabling LLMs to compensate for their inherent limitations. For example, grounding and verification mechanisms help reduce hallucination, controlled execution improves reliability, and explicit role decomposition mitigates long-context instability. Agent systems thus do not merely wrap LLMs—they provide the organizational scaffolding needed to transform a general-purpose model into a dependable data science practitioner.
Seen from this perspective, our analysis focuses on four design dimensions: agent role design (§3.1), execution structure (§3.2), external knowledge integration (§3.3), and reflection mechanisms (§3.4). They form a structured solution to the inherent limitations of LLMs in data science settings. The remainder of this section examines each dimension in turn, showing how existing systems operationalize these design principles in practice.
@@ 3.1.1 Single Agent
The single-agent design is the simplest design structure as shown in Figure 4. Most existing single-agent works directly follow the ReAct design (a prompting technique , where LLMs are used to generate both reasoning traces and task-specific actions) freimanis2024tableanalyst; chen2024multi; sun2024lambda; jing2024dsbench; hu2024infiagent; deng2024pentestgpt; le2023codechain; zhang2024codeagent; gupta2024codenavbendinelli2025exploring; xu2025dagent, where the single agent will perform thought, action, and observation processes iteratively on its own.
Single-agent designs offer minimal overhead and are easy to deploy, making them attractive for low-latency or low-cost settings. However, since all reasoning, memory, and execution are concentrated in a single model, they inherit most of the LLM limitations: hallucinations are unmitigated, generated code is fragile, and long-horizon tasks often suffer from goal drift and lost context. As a result, single-agent systems are often insufficient for complex or multi-stage workflows that require long-horizon reasoning, consistent state tracking, or robust code execution.
@@ 3.1.2 Two Agents
Planner and executor. Different from the single agent discussed previously, some works split the single agent into two agents huang2024foodpuzzle; liu2024largehyper; zhang2023datacopilot; chi2024sela; li2024tapilotwang2025dsmentor, planner and executor. As an example shown in Figure 5, the planner will get observations from previous execution results or users’ requests, and generate a next-step plan, or a whole plan in advance. Then, the executor is prompted to interact with the environments and will follow the generated plans to execute.
@@
Coder and reviewer. Another type of two-agent structure is the “coder and reviewer” style trirat2024automl; huang2023benchmarking as shown in Figure 6. In such a design, the coder is responsible for completing tasks following the typical ReAct structure. Another agent, the reviewer is introduced to check the validity of the code generated by the coder. trirat2024automl allows the reviewer to check the generated response at each step of generation, while huang2023benchmarking only allows the reviewer to function at the end of the entire generation process.
Two-agent designs improve over single-agent systems by splitting responsibilities into complementary roles. Planner–executor structures provide more structured task decomposition, and coder–reviewer structures introduce lightweight verification, catching some errors or hallucinations that a single agent would miss. However, the approach remains sensitive to failures such as a planner hallucinating an incorrect plan or a reviewer failing to identify subtle logical errors. Communication cost also increases, and the reliability gains depend heavily on how well the two roles align. As a result, two-agent systems offer a moderate balance between simplicity and robustness, but their effectiveness is bounded by the underlying LLM’s consistency and self-correction ability.
@@ 3.1.3 Multiple Agents
Software engineering-style team. The Software Engineering (SE) team-style agent design draws inspiration from traditional human software development teams qian2024chatdev; zhao2024visioncoder; trirat2024automl; hong2023metagpt; tao2024magis; lin2024soen101; nguyen2024agilecoder. In an SE team-style framework, agents are typically assigned roles that correspond to key human roles in software development, here we state some example roles: The product manager defines the product vision and organizes requirements. The requirements analyst translates user needs into detailed software specifications. The scrum master facilitates task planning and sprint coordination. Hierarchical roles like team leader, module Leader, and function coordinator handle task decomposition at varying levels of granularity. The executor for the task is the developer, who is responsible for implementing the code, while the senior developer refines it. Finally, the QA engineer ensures quality through rigorous testing, and the tester validates specific functionalities. Some SE team-style agents also perform hierarchical roles which allow some agents with high permissions to manage others.
@@
Minimum function agents. Minimum function agents are designed to handle narrowly scoped and atomic tasks separately, as shown in Figure 7.
Across existing frameworks, minimum function agents are given very small and specific functions to handle. For instance, code search agents locate relevant files, classes, or methods within a repository, while fault localization agents identify buggy code sections using debugging techniques like spectrum-based fault localization zhang2024autocoderover. Other agents specialize in generating patches to fix identified issues, executing tests to validate code correctness, or systematically building repository structures from high-level descriptions zan2024codes; arora2024masai. In the context of exception handling, some agents detect fragile code, identify exception types, and implement robust handling mechanisms to enhance code reliability zhang2024seeker. All works emphasize task decomposition and modularity, with outputs from one agent often serving as inputs for another, forming structured and collaborative workflows zan2024codes; phan2024hyperagent; arora2024masaiseo2025spio; you2025datawiseagent; ou2025automind.
@@
Client-server design. Client-server agents, adopt a hierarchical architecture where a central controller agent manages and coordinates the operations of multiple specialized client agents zhang2023automl; yang2024finrobot; gandhi2024budgetmlagent; zhao2024lightva.
The controller or server agent functions as a project manager, planning entire workflows of the given tasks, and allocates the tasks to the client agents. Client agents perform as roles like software engineers or testers, focusing on executing specific subtasks such as data analysis, modeling, or feedback generation. These roles allow for clear separation of tasks, while also enabling dynamic adjustments based on task requirements or external feedback zhang2023automl; bai2024multi; zhao2024lightva; shen2024data.
Multi-agent architectures offer the highest scalability and strongest resistance to long-context forgetting, as each agent operates within a narrowly scoped role. They can achieve high reliability through division of labor, modular workflows, and cross-agent verification. Yet the increased number of agents introduces coordination overhead, susceptibility to miscommunication, and longer execution latency. These systems are powerful but harder to control, and their effectiveness depends on careful manual workflow design.
@@ 3.1.4 Dynamic Agents
Dynamic agents represent a class of multi-agent systems that emphasize adaptability by dynamically creating, modifying, or expanding agents during runtime. Unlike static agents, which follow predefined workflows and configurations, dynamic agents are designed to respond to the complexity or variability of tasks by adjusting the agents’ internal structure (prompt, etc.) or introducing new agents hu2024self; ishibashi2024self. Current frameworks for dynamic agent creation adopt two primary paradigms:
@@
Hierarchical agent generation. This paradigm involves a parent agent or a high-level controller (e.g., Mother Agent in SoA ishibashi2024self) that decomposes complex tasks into subtasks and creates child agents to handle the subtasks. Each child agent operates independently on its specific subtask. It is particularly effective in managing tasks with clear functional divisions, such as modular code generation or system-level software design ishibashi2024self.
@@
Iterative agent generation through feedback. Dynamic agents adjust their structures and behaviors iteratively based on real-time feedback from their environment or other agents (Figure 10). EvoMAC hu2024self exemplifies this paradigm, employing a collaborative rather than hierarchical approach, where agents refine their outputs and workflows through mechanisms analogous to backpropagation—such as a loss agent computing errors and an update agent adjusting agent workflows accordingly. This textual feedback enables the dynamic creation or reconfiguration of agents, supporting adaptability for tasks with evolving requirements, such as changing software specifications (Figure 10) hu2024self.
Dynamic agent generation offers high flexibility by creating agents on demand rather than relying on manually designed roles, prompts, and pipeline. This allows the system to adjust its structure at runtime and better accommodate diverse or evolving data science tasks. However, dynamically spawning agents also amplifies LLM uncertainty, reduces reproducibility, and increases computational and coordination overhead. As a result, these systems provide strong adaptability but exhibit lower predictability and stability, making them challenging to deploy in stable production settings.
@@ 3.1.5 Summary of Agent Role
    With a manager: In this structure, a central agent manages and controls all agents. Software engineering-style agents and Client-server agents mainly have this structure. For example, in the AutoML-GPTtrirat2024automl framework, a central LLM serves as the controller, managing the entire pipeline by integrating specialized agents for subtasks such as model design and hyperparameter tuning.
@@
    Without a manager: Each agent operates independently and solves tasks autonomously. Minimum function agents mainly pose this structure, since all the agents with minimum function share the same position. An example of this is the MASAI arora2024masai framework, which utilizes decentralized agents that collaborate on machine learning and data science tasks by sharing results but not a central management system.
@@
    Hierarchical managers: A higher-level agent controls lower-level agents in a layered structure. Hierarchically generated agents and part of software engineering style agents mainly pose such structure. An example of this can be found in Hierarchical agent generation, such as in EvoMAC hu2024self, where a parent agent dynamically creates child agents to handle specific subtasks during runtime.
Dimension Single Agent Two-Agent Multi-Agent Dynamic Agents
Reliability
Scalability
Coordination Cost
Predictability / Stability
Industrial Applicability Quick analyses, ad-hoc queries, and low-risk automation Medium-scale workflows requiring basic verification and predictable execution Production pipelines with modular stages, quality checks, and stable data dependencies Exploratory analytics, quick pipeline prototyping, or changeable environment
Table 3: Trade-off summary across agent role designs. ●= strong,  ◐= moderate,  ○= weak.
@@
    Compete: Agents work against each other to complete a task, often in the coder-reviewer paradigm (similar to the adversarial concept in GAN). In frameworks like MASAI arora2024masai, agents engage in competitive strategies to address machine learning and data science challenges. The competition between reviewers and coders allows the iterative refinement of the code. Some works also allow multiple agents to propose multiple plans to compete.
@@
    Collaborate: Agents work together toward a shared objective. For example, in the MAGIS tao2024magis framework, agents assume different roles like Manager, Developer, and QA Engineer to collaborate on resolving GitHub issues. Their tasks are divided to ensure modular development, with continuous collaboration between agents.
@@
    Hybrid:
    Agents alternate between competing and collaborating based on the task requirements. For instance, in AutoCodeRover zhang2024autocoderover, agents work together to localize faults and generate patches, but may compete in terms of optimizing solutions or strategies based on the specific issue at hand.
@@
    Static Task Allocation: In some systems, agents are assigned a fixed set of tasks that they perform. For example, in the Data Director hong2024data framework, agents follow a static task allocation, where the tasks are predefined and agents work through structured stepwise execution.
@@
    Dynamic Task Allocation: Tasks are allocated based on the real-time needs and feedback from the system or environment. An example of dynamic task allocation can be seen in the EvoMAC hu2024self framework, where agents adjust dynamically based on environmental feedback, creating or dismissing agents as needed to refine or expand their tasks.
@@
    Coarse Granularity: Some agents are given broader, less detailed tasks. For example, in AutoML-GPT trirat2024automl, the central controller agent coordinates the entire machine learning pipeline, handling coarse-grained tasks such as managing the overall workflow rather than focusing on the details of each individual task.
@@
    Fine Granularity: In other cases, tasks are broken down into smaller units for more specific execution. For example, in MapCoder islam2024mapcoder, multiple agents collaborate, each handling a fine-grained task such as code generation, debugging, or retrieval. This detailed task assignment ensures high accuracy but requires more computation overhead.